OpenStudy (anonymous):

Admission to the zoo costs $4 for adults and $2 for kids. The total bill for your entire extended family was $30. This scenario can be modeled by the equation 4a + 2k = 30. Which is not a possible combination of adults and kids? 5 adults and 5 kids 4 adults and 7 kids 7 adults and 1 kid 3 adults and 6 kids

OpenStudy (anonymous):

like guess and check in this situation. 5(4$) + 5*2 = $30 4($4) + 7*($2) = $30 7($4) + 1*($2) = $30 3*($4) + 6($2) = $24 3 adults and 6 kids is not possible as it would only cost $24 to get in.